OpenStudy (anonymous):

\[\sqrt{2x -3} - \sqrt{x+7} = 4\]

hartnn (hartnn):

square both sides.

hartnn (hartnn):

what u get ?

OpenStudy (anonymous):

Take one sqrt term to right that would be easy

hartnn (hartnn):

i doubt that... just the same thing....squaring twice will be necessary

OpenStudy (anonymous):

wait 1 min plz

OpenStudy (anonymous):

i dont understand the 3rd setp

hartnn (hartnn):

its squared both sides in third step

hartnn (hartnn):

\((x-26)^2=8^2\sqrt{x+7}^2\)

hartnn (hartnn):

for left side use \((x-a)^2=x^2-2ax+a^2\)

OpenStudy (anonymous):

okk

hartnn (hartnn):

any more doubts ?

OpenStudy (anonymous):

yes i got it:)

OpenStudy (anonymous):

nop

hartnn (hartnn):

good...

OpenStudy (anonymous):

i have doubt one question

hartnn (hartnn):

which >

OpenStudy (anonymous):

this one

OpenStudy (anonymous):

the question is: \[x=2y\] \[\sqrt{x+y} - \sqrt{x-y} = 2\]

hartnn (hartnn):

which step ?

OpenStudy (anonymous):

All setps

hartnn (hartnn):

first you replace x by 2y then square both sides.

hartnn (hartnn):

what u get ?

hartnn (hartnn):

isn't it x=2y+1 ??

OpenStudy (anonymous):

is system equation